# LumiMAS: A Comprehensive Framework for Real-Time Monitoring and Enhanced Observability in Multi-Agent Systems

### v1

@@ C.3 DPI Attack Evaluation
Application
DPI
Scenario
Method Performance Overhead
Acurecy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Cost $\downarrow$
Genfic Misinformation LLM-as-a-judge 0.870 0.885 1.000 0.794 0.260 17.738 7910.300
Agent-as-a-judge 0.930 0.935 1.000 0.877 0.140 23.886 5772818.275
@@
TripPlaner Misinformation LLM-as-a-judge 0.950 0.952 1.000 0.909 0.100 12.402 6615.430
Agent-as-a-judge 0.920 0.926 1.000 0.862 0.160 21.967 3541434.755
@@
TripPlaner Exhaustion LLM-as-a-judge 0.960 0.962 1.000 0.926 0.080 11.949 5943.520
Agent-as-a-judge 0.925 0.930 0.990 0.876 0.140 14.514 3055570.145
@@
TripPlaner Backdoor LLM-as-a-judge 0.915 0.915 0.920 0.911 0.090 12.363 6742.690
Agent-as-a-judge 0.885 0.883 0.870 0.897 0.100 14.985 3108083.005
@@
Table 9: Results on LangGraph applications with GPT-4o-mini as the underlying model.
DPI
Scenario
Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Cost $\downarrow$
Hallucination Consistency of 3 0.655 0.693 0.780 0.624 0.470 35.726 7836.735
Consistency of 5 0.595 0.649 0.750 0.573 0.560 61.103 13006.280
LLM-as-a-judge 0.585 0.303 0.180 0.947 0.010 9.307 2468.660
Agent-as-a-judge 0.550 0.237 0.140 0.778 0.040 11.463 1322839.065
@@
LLM-as-a-judge 0.785 0.768 0.710 0.835 0.140 7.614 2266.500
Agent-as-a-judge 0.840 0.822 0.740 0.925 0.060 9.773 1244947.160
@@
DPI
Misinformation
LLM-as-a-judge 0.840 0.862 1.000 0.758 0.320 11.407 6024.035
Agent-as-a-judge 0.840 0.862 1.000 0.758 0.320 22.888 4637629.795
EPI (Ours) 0.785 0.823 1.000 0.699 0.430 0.0046 -
@@
DPI
Exhaustion
LLM-as-a-judge 0.770 0.807 0.960 0.696 0.42 12.933 6044.565
Agent-as-a-judge 0.825 0.848 0.980 0.748 0.33 20.094 4319530.840
@@
DPI
Backdoor
LLM-as-a-judge 0.745 0.756 0.790 0.725 0.3 12.752 6284.210
Agent-as-a-judge 0.735 0.731 0.720 0.742 0.25 20.3111 4011112.610
@@
Memory
Poisoning
LLM-as-a-judge 0.830 0.830 0.830 0.830 0.170 10.032 3679.995
Agent-as-a-judge 0.715 0.671 0.580 0.795 0.150 15.906 1722696.895
EPI (Ours) 0.585 0.618 0.670 0.573 0.670 0.0029 -
@@
Table 10: Results on CrewAI applications with o3-mini as the underlying model.
Results on o3-mini.
In this subsection, we explore our method using a different underlying model to demonstrate its effectiveness and robustness. Table 10 shows the effectiveness of our anomaly detection methods on the CrewAI applications using o3-mini as the underlying model. The table presents the performance across four evaluated vulnerability types (rows), along with each detection approach’s performance metrics and overhead. we observed that logs generated with o3-mini as the underlying LLM are relatively shorter than those generated with GPT-4o-mini, as the o3-mini model produces shorter agent-LLM interaction cycles, probably due to the high reasoning capabilities of the LLM. our approach is the most efficient in terms of runtime and has zero inference cost regarding token usage, in contrast to alternative baselines that rely on external API calls. Our method achieved strong results in detecting DPI attacks, as shown in the results table. However, the combined approach is relatively less effective at detecting hallucination in the o3-mini model logs. This may be due to the shorter reasoning cycles causing a reduced amount of textual data, which leads to logs exhibiting less distinguishable patterns. In the MP scenario, the Semantic and Combined detection approaches, as well as the LLM-based baselines, perform reasonably well in contrast to the EPI approach due to the textual nature of the attack. Interestingly, in the case of bias detection, the o3-mini logs exhibited more distinct patterns, and the EPI-based approach achieved strong results in contrast to the semantic approach.
@@
Figure 8: RCA results obtained using the CrewAI apps, with GPT-o3-mini as the underlying model.
Figure 9: Anomaly classification performance confusion matrix with GPT-o3-mini as the underlying model.
Anomaly Detection Results On LangGraph Applications
In this subsection, we evaluate our method on a different MAS platform to demonstrate its robustness. Table 9 presents the LangGraph framework’s evaluation of 3 different DPI attacks (Misinformation, Exhaustion, and Backdoor) on the Trip Planner application as well as the misinformation DPI attack on the Genfic application, as outlined in the first and second columns. The results indicate that our combined approach achieves the best performance on the Genfic application, attaining near-perfect scores across all evaluation metrics. Although the baselines achieved strong results on the Trip Planner application, their high cost and inference time make them impractical for real-time detection. In contrast, our semantic approach delivers competitive results while maintaining a low average inference time and zero cost, making it more suitable for real-time detection.
Larger Scale Scenario Time Consumption
In this subsection, we conducted a larger-scale experiment to demonstrate our anomaly detection method’s efficiency, even when applied to larger application settings. In the examined case, we processed logs containing nearly double the number of agents (10), resulting in a proportional increase in log size. The measured average inference times (across 5 experiments) were 0.008, 0.150, and 0.157 seconds for EPI, Semantic, and Combine, respectively. The observed inference time remains substantially low and still enables real-time anomaly detection.
Enhanced Anomaly Detection Result Using the Explanation Layer
@@
Failure Method
With
CLS
Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Cost $\downarrow$
Hallucination EPI No 0.745 0.730 0.690 0.775 0.200 0.003 -
Embeddings No 0.720 0.682 0.600 0.790 0.160 0.014 -
Combine No 0.765 0.728 0.630 0.863 0.100 0.017 -
EPI Yes 0.770 0.709 0.560 0.965 0.020 7.764 6715.610
Embeddings Yes 0.770 0.705 0.550 0.9821 0.010 6.365 6042.800
Combine Yes 0.780 0.725 0.580 0.9667 0.020 6.555 5584.355
Bias EPI No 0.625 0.675 0.780 0.595 0.530 0.016 -
Embeddings No 0.600 0.615 0.640 0.593 0.440 0.020 -
Combine No 0.660 0.670 0.690 0.651 0.370 0.024 -
EPI Yes 0.635 0.664 0.720 0.6154 0.450 10.2638 13621.970
Embeddings Yes 0.595 0.589 0.580 0.5979 0.390 8.4898 9918.080
Combine Yes 0.655 0.6425 0.620 0.667 0.310 8.905 11527.800
Direct
Prompt
Injection (DPI)
Misinformation
EPI No 0.795 0.827 0.980 0.715 0.390 0.008 -
Embeddings No 0.815 0.843 0.990 0.733 0.360 0.071 -
Combine No 0.815 0.830 0.90 0.769 0.270 0.084 -
EPI Yes 0.910 0.906 0.870 0.946 0.050 14.314 50466.635
Embeddings Yes 0.880 0.875 0.840 0.913 0.080 13.643 49090.870
Combine Yes 0.840 0.824 0.750 0.915 0.070 12.744 43255.985
Direct
Prompt
Injection (DPI)
Exhaustion
EPI no 0.805 0.84 1.000 0.719 0.390 0.009 -
Embeddings No 0.800 0.828 0.960 0.727 0.360 0.082 -
Combine No 0.865 0.881 1.000 0.787 0.270 0.096 -
EPI Yes 0.915 0.913 0.890 0.937 0.060 13.468 45144.945
Embeddings Yes 0.880 0.878 0.860 0.896 0.10 12.880 43381.245
Combine Yes 0.895 0.888 0.830 0.954 0.04 12.152 40430.070
Direct
Prompt
Injection (DPI)
Backdoor
EPI No 0.625 0.630 0.640 0.621 0.390 0.007 -
Embeddings No 0.695 0.711 0.750 0.676 0.360 0.067 -
Combine No 0.620 0.573 0.510 0.654 0.270 0.083 -
EPI Yes 0.755 0.710 0.600 0.870 0.090 9.153 36858.510
Embeddings Yes 0.825 0.807 0.730 0.901 0.080 10.390 40297.325
Combine Yes 0.730 0.649 0.500 0.926 0.040 7.495 28811.960
Indirect
Prompt
Injection (IPI)
EPI No 0.965 0.966 1.000 0.935 0.070 0.029 -
Embeddings No 0.890 0.894 0.930 0.861 0.150 0.102 -
Combine No 0.970 0.970 0.970 0.970 0.030 0.142 -
EPI Yes 0.940 0.936 0.880 1.000 0.000 11.088 37395.645
Embeddings Yes 0.910 0.902 0.830 0.988 0.010 10.966 37641.450
Combine Yes 0.925 0.919 0.850 1.000 0.000 10.819 34951.155
Memory
Poisoning (MP)
EPI No 0.540 0.643 0.830 0.525 0.750 0.003 -
Embeddings No 0.690 0.710 0.760 0.667 0.380 0.0283 -
Combine No 0.500 0.561 0.640 0.500 0.640 0.0315 -
EPI Yes 0.770 0.760 0.730 0.794 0.190 18.893 28162.810
Embeddings Yes 0.805 0.785 0.710 0.877 0.100 10.713 20317.920
Combine Yes 0.690 0.659 0.600 0.732 0.220 12.281 26117.735
Table 11: Anomaly detection with and without classification Results compared on all applications with GPT-4o-mini as the underlying model on CrewAI applications.
Vulnerability Method
With
CLS
Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$ Cost $\downarrow$
Genfic EPI No 0.980 0.980 0.970 0.990 0.010 0.007 -
Embeddings No 0.955 0.955 0.95 0.960 0.040 0.049 -
Combine No 0.990 0.990 0.990 0.990 0.010 0.052 -
EPI Yes 0.980 0.980 0.960 1.000 0.000 11.041 18052.885
Embeddings Yes 0.945 0.944 0.930 0.959 0.040 12.274 18402.970
Combine Yes 0.985 0.985 0.980 0.990 0.010 10.232 18404.170
Direct
Prompt
Injection (DPI)
Misinformation
EPI No 0.610 0.625 0.650 0.602 0.430 0.008 -
Embeddings No 0.895 0.893 0.880 0.907 0.090 0.076 -
Combine No 0.680 0.698 0.740 0.661 0.380 0.086 -
EPI Yes 0.715 0.637 0.500 0.877 0.070 10.173 48256.435
Embeddings Yes 0.845 0.817 0.690 1.000 0.000 9.058 43967.960
Combine Yes 0.815 0.793 0.710 0.899 0.080 12.546 50288.775
Direct
Prompt
Injection (DPI)
Exhaustion
EPI No 0.785 0.823 1.000 0.699 0.430 0.0214 -
Embeddings No 0.955 0.957 1.000 0.916 0.090 0.288 -
Combine No 0.810 0.840 1.000 0.725 0.380 0.296 -
EPI Yes 0.930 0.932 0.960 0.906 0.100 14.264 57154.240
Embeddings Yes 0.975 0.975 0.980 0.970 0.030 10.985 41082.205
Combine Yes 0.945 0.945 0.940 0.940 0.050 13.705 53576.620
Direct
Prompt
Injection (DPI)
Backdoor
EPI No 0.570 0.570 0.570 0.570 0.430 0.009 -
Embeddings No 0.890 0.888 0.870 0.906 0.090 0.081 -
Combine No 0.690 0.710 0.790 0.667 0.380 0.088 -
EPI Yes 0.740 0.680 0.550 0.887 0.070 10.927 42188.810
Embeddings Yes 0.905 0.898 0.840 0.966 0.030 10.219 41285.840
Combine Yes 0.815 0.793 0.710 0.899 0.080 12.546 50288.775
Table 12: Anomaly detection with and with hout classification Results compared on all applications with GPT-4o-mini as the underlying model on Langgraph.
Anomaly Explanation Evaluation
@@
  • Explanation Completeness (0–3 points): Evaluates whether the explanation includes both the classification decision and the predicted root cause, forming a coherent and comprehensive rationale.
Table 13 presents the evaluation scores across the different anomaly types. Notably, pattern detection scores were higher for anomaly types involving explicit manipulations, such as DPI, IPI, and MP, compared to the more implicit categories of Bias and Hallucination. In particular, the DPI cases consistently achieved strong scores across all dimensions, with a pattern detection score of 3.5, evidence quality of 2.4, and completeness of 2.2. IPI cases showed similarly balanced performance. In contrast, hallucination and bias explanations were less consistent, reflecting the difficulty in articulating abstract or ambiguous failure modes. These types scored lower in both pattern detection and completeness, suggesting a need for further refinement when handling subtle semantic failures. The results show an overall average score of 7.5 out of 10, indicating that the generated explanations were generally reliable and informative. To further validate these findings, we conducted a human-in-the-loop review. A subset of explanations was manually assessed to verify their semantic relevance and consistency with the known failure conditions. This review confirmed that the generated explanations aligned with expectations and effectively supported the interpretation of the detected anomaly.
@@
Table 13: Anomaly explanation evaluation scores across anomaly types with GPT-4o-mini as the underlying model on CrewAI applications.

### v2

@@ C.2.7. Unsupervised Classification
Table 11. Results on LangGraph applications with GPT-4o-mini as the underlying model
Application DPI Scenario Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
Genfic Misinformation LLM-as-a-judge 0.870 0.885 1.000 0.794 0.260 17.738 7910.3
Agent-as-a-judge 0.930 0.935 1.000 0.877 0.140 23.886 5772818.3
LogBERT 0.500 0.667 1.000 0.500 1.000 0.003 -
@@
TripPlanner Misinformation LLM-as-a-judge 0.950 0.952 1.000 0.909 0.100 12.402 6615.4
Agent-as-a-judge 0.920 0.926 1.000 0.862 0.160 21.967 3541434.8
LogBERT 0.51 0.671 1.000 0.505 0.980 0.039 -
@@
TripPlanner Exhaustion LLM-as-a-judge 0.960 0.962 1.000 0.926 0.080 11.949 5943.5
Agent-as-a-judge 0.925 0.930 0.990 0.876 0.140 14.514 3055570.2
LogBERT 0.510 0.671 1.000 0.505 0.980 0.047 -
@@
TripPlanner Backdoor LLM-as-a-judge 0.915 0.915 0.920 0.911 0.090 12.363 6742.7
Agent-as-a-judge 0.885 0.883 0.870 0.897 0.100 14.985 3108083.0
LogBERT 0.460 0.625 0.900 0.479 0.980 0.024 -
@@
Table 12. Results on CrewAI applications with o3-mini as the underlying model
Vulnerability Method Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
Hallucination Consistency of 3 0.655 0.693 0.780 0.624 0.470 35.726 7836.7
Consistency of 5 0.595 0.649 0.750 0.573 0.560 61.103 13006.3
LLM-as-a-judge 0.585 0.303 0.180 0.947 0.010 9.307 2468.7
Agent-as-a-judge 0.550 0.237 0.140 0.778 0.040 11.463 1322839.1
LogBERT 0.500 0.667 1.000 0.500 1.000 0.037 -
@@
LLM-as-a-judge 0.785 0.768 0.710 0.835 0.140 7.614 2266.5
Agent-as-a-judge 0.840 0.822 0.740 0.925 0.060 9.773 1244947.2
LogBERT 0.545 0.209 0.120 0.800 0.03 0.027 -
@@
DPI Misinformation LLM-as-a-judge 0.840 0.862 1.000 0.758 0.320 11.407 6024.0
Agent-as-a-judge 0.840 0.862 1.000 0.758 0.320 22.888 4637629.8
LogBERT 0.555 0.491 0.430 0.573 0.320 0.034 -
EPI (Ours) 0.785 0.823 1.000 0.699 0.430 0.005 -
@@
DPI Exhaustion LLM-as-a-judge 0.770 0.807 0.960 0.696 0.420 12.933 6044.6
Agent-as-a-judge 0.825 0.848 0.980 0.748 0.330 20.094 4319530.8
LogBERT 0.840 0.862 1.000 0.757 0.320 0.016 -
@@
DPI Backdoor LLM-as-a-judge 0.745 0.756 0.790 0.725 0.30 12.752 6284.2
Agent-as-a-judge 0.735 0.731 0.720 0.742 0.250 20.3111 4011112.6
LogBERT 0.610 0.580 0.540 0.628 0.320 0.014 -
@@
MP LLM-as-a-judge 0.830 0.830 0.830 0.830 0.170 10.032 3680.0
Agent-as-a-judge 0.715 0.671 0.580 0.795 0.150 15.906 1722696.9
LogBERT 0.765 0.768 0.780 0.757 0.250 0.009 -
EPI (Ours) 0.585 0.618 0.670 0.573 0.670 0.003 -
@@
C.2.8. Results on o3-mini
In this subsection, we explore our method using a different underlying model to demonstrate its effectiveness and robustness. Table 12 shows the effectiveness of our anomaly detection methods on the CrewAI applications using o3-mini as the underlying model. The table presents the performance across four evaluated vulnerability types (rows), along with each detection approach’s performance metrics and overhead. we observed that logs generated with o3-mini as the underlying LLM are relatively shorter than those generated with GPT-4o-mini, as the o3-mini model produces shorter agent-LLM interaction cycles, probably due to the high reasoning capabilities of the LLM. our approach is the most efficient in terms of runtime and has zero inference cost regarding token usage, in contrast to alternative baselines that rely on external API calls. Our method achieved strong results in detecting DPI attacks, as shown in the results table. However, the combined approach is relatively less effective at detecting hallucination in the o3-mini model logs. This may be due to the shorter reasoning cycles causing a reduced amount of textual data, which leads to logs exhibiting less distinguishable patterns. In the MP scenario, the Semantic and Combined detection approaches, as well as the LLM-based baselines, perform reasonably well in contrast to the EPI approach due to the textual nature of the attack. Interestingly, in the case of bias detection, the o3-mini logs exhibited more distinct patterns, and the EPI-based approach achieved strong results in contrast to the semantic approach. While LogBERT offers lower runtime than LumiMAS’s combined approach, its overall performance is notably weaker in most cases. It excels when anomalies manifest through structural deviations in the log sequence (e.g., the DPI Exhaustion scenario), effectively leveraging sequential patterns to identify such irregularities. In contrast, LogBERT struggles with failures that are not structurally evident but instead stem from semantic inconsistencies or performance-related issues at the system level (e.g., DPI hallucination and bias).
@@ C.2.8. Results on o3-mini
Figure 8. Root cause analysis (RCA) results obtained using CrewAI apps with GPT-o3-mini as the underlying model
[…]
Root cause analysis (RCA) results obtained using CrewAI apps, across four vulnerability types with GPT-o3-mini as the underlying model. The figure compares the proposed RCA agent with baseline methods, with grouped bars representing each method for each vulnerability category. The vertical axis indicates performance scores, and the horizontal axis lists the vulnerability types.
Figure 9. Anomaly classification performance confusion matrix on CrewAI apps with GPT-o3-mini as the underlying model.
[…]
Confusion matrix showing anomaly classification performance on on CrewAI apps with GPT-o3-mini as the underlying model. Rows correspond to true classes and columns correspond to predicted classes. Cell values indicate the number of predictions for each class combination, and color intensity reflects the magnitude of the counts
C.2.9. Anomaly Detection Results On LangGraph Applications
In this subsection, we evaluate our method on a different MAS platform to demonstrate its robustness. Table 11 presents the LangGraph framework’s evaluation of 3 different DPI attacks (Misinformation, Exhaustion, and Backdoor) on the Trip Planner application as well as the misinformation DPI attack on the Genfic application, as outlined in the first and second columns. The results indicate that our combined approach achieves the best performance on the Genfic application, attaining near-perfect scores across all evaluation metrics. Although the LLM-based baselines achieved strong results on the Trip Planner application, their high cost and inference time make them impractical for real-time detection. In contrast, our semantic approach delivers competitive results while maintaining a low average inference time and zero cost, making it more suitable for real-time detection. LogBERT, as a sequential approach, did not perform well on the DPI attacks, since the LangGraph logs did not exhibit significant structural changes. As a result, it failed to effectively distinguish between benign and anomalous logs.
C.2.10. Larger Scale Scenario Time Consumption
In this subsection, we conducted a larger-scale experiment to demonstrate our anomaly detection method’s efficiency, even when applied to larger application settings. In the examined case, we processed logs containing nearly double the number of agents (10), resulting in a proportional increase in log size. The measured average inference times (across 5 experiments) were 0.008, 0.150, and 0.157 seconds for EPI, Semantic, and Combined, respectively. The observed inference time remains substantially low and still enables real-time anomaly detection.
C.2.11. Enhanced Anomaly Detection Result Using the Explanation Layer
@@
C.2.12. Anomaly Explanation Evaluation
@@
  • Explanation Completeness (0–3 points): Evaluates if the explanation includes both the classification decision and the predicted root cause, forming a coherent and comprehensive rationale.
Table 13 presents the evaluation scores across the different anomaly types. Notably, pattern detection scores were higher for anomaly types involving explicit manipulations, such as DPI, IPI, and MP, compared to the more implicit categories of Bias and Hallucination.
In particular, the DPI cases consistently achieved strong scores across all dimensions, with a pattern detection score of 3.5, evidence quality of 2.4, and completeness of 2.2. IPI cases showed similarly balanced performance. In contrast, hallucination and bias explanations were less consistent, reflecting the difficulty in articulating abstract or ambiguous failure modes. The results show an overall average score of 7.5 out of 10, indicating that the generated explanations were generally reliable and informative.
To further validate our findings, we conducted a human evaluation of the anomaly explanations. Following an application-grounded evaluation doshi2017towards, three domain experts independently rated a subset of explanations for each failure type using a similar structured rubric like the automated evaluation with three dimensions: Pattern Detection (0–4), Evidence Quality (0–3), and Completeness (0–3). Each expert received the log snippet, the system-generated explanation, and the corresponding ground-truth failure label. Table 16 reports the mean expert scores for each dimension and failure type. Bias-related explanations achieved the highest quality ratings, while DPI explanations were the most challenging for experts to interpret. Overall, the explanations received a mean score of 7.3 out of 10, indicating that they were relevant and coherent. The experts’ review confirmed that the generated explanations aligned with the expected failure characteristics and effectively supported the interpretation of the detected anomalies.
Table 13. Anomaly explanation evaluation scores across anomaly types with GPT-4o-mini as the underlying model on CrewAI applications.
@@
Table 14. Anomaly detection with and without classification Results compared on all applications with GPT-4o-mini as the underlying model on CrewAI applications.
Failure Method With CLS Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
Hallucination EPI No 0.745 0.730 0.690 0.775 0.200 0.003 -
Embeddings No 0.720 0.682 0.600 0.790 0.160 0.014 -
Combined No 0.765 0.728 0.630 0.863 0.100 0.017 -
EPI Yes 0.770 0.709 0.560 0.965 0.020 7.764 6715.6
Embeddings Yes 0.770 0.705 0.550 0.982 0.010 6.365 6042.8
Combined Yes 0.780 0.725 0.580 0.967 0.020 6.555 5584.4
Bias EPI No 0.625 0.675 0.780 0.595 0.530 0.016 -
Embeddings No 0.600 0.615 0.640 0.593 0.440 0.020 -
Combined No 0.660 0.670 0.690 0.651 0.370 0.024 -
EPI Yes 0.635 0.664 0.720 0.615 0.450 10.264 13621.9
Embeddings Yes 0.595 0.589 0.580 0.598 0.390 8.490 9918.1
Combined Yes 0.655 0.643 0.620 0.667 0.310 8.905 11527.8
DPI Misinformation EPI No 0.795 0.827 0.980 0.715 0.390 0.008 -
Embeddings No 0.815 0.843 0.990 0.733 0.360 0.071 -
Combined No 0.815 0.830 0.90 0.769 0.270 0.084 -
EPI Yes 0.910 0.906 0.870 0.946 0.050 14.314 50466.6
Embeddings Yes 0.880 0.875 0.840 0.913 0.080 13.643 49090.8
Combine Yes 0.840 0.824 0.750 0.915 0.070 12.744 43255.9
DPI Exhaustion EPI No 0.805 0.84 1.000 0.719 0.390 0.009 -
Embeddings No 0.800 0.828 0.960 0.727 0.360 0.082 -
Combined No 0.865 0.881 1.000 0.787 0.270 0.096 -
EPI Yes 0.915 0.913 0.890 0.937 0.060 13.468 45144.9
Embeddings Yes 0.880 0.878 0.860 0.896 0.100 12.880 43381.2
Combined Yes 0.895 0.888 0.830 0.954 0.040 12.152 40430.1
DPI Backdoor EPI No 0.625 0.630 0.640 0.621 0.390 0.007 -
Embeddings No 0.695 0.711 0.750 0.676 0.360 0.067 -
Combined No 0.620 0.573 0.510 0.654 0.270 0.083 -
EPI Yes 0.755 0.710 0.600 0.870 0.090 9.153 36858.5
Embeddings Yes 0.825 0.807 0.730 0.901 0.080 10.390 40297.3
Combined Yes 0.730 0.649 0.500 0.926 0.040 7.495 28811.9
IPI EPI No 0.965 0.966 1.000 0.935 0.070 0.029 -
Embeddings No 0.890 0.894 0.930 0.861 0.150 0.102 -
Combined No 0.970 0.970 0.970 0.970 0.030 0.142 -
EPI Yes 0.940 0.936 0.880 1.000 0.000 11.088 37395.6
Embeddings Yes 0.910 0.902 0.830 0.988 0.010 10.966 37641.5
Combined Yes 0.925 0.919 0.850 1.000 0.000 10.819 34951.1
MP EPI No 0.540 0.643 0.830 0.525 0.750 0.003 -
Embeddings No 0.690 0.710 0.760 0.667 0.380 0.0283 -
Combined No 0.500 0.561 0.640 0.500 0.640 0.0315 -
EPI Yes 0.770 0.760 0.730 0.794 0.190 18.893 28162.8
Embeddings Yes 0.805 0.785 0.710 0.877 0.100 10.713 20317.9
Combined Yes 0.690 0.659 0.600 0.732 0.220 12.281 26117.7
Table 15. Anomaly detection with and without classification Results compared on all applications with GPT-4o-mini as the underlying model on Langgraph.
Vulnerability Method With CLS Performance Overhead
Accuracy $\uparrow$ F1 $\uparrow$ Recall $\uparrow$ Precision $\uparrow$ FPR $\downarrow$ Latency $\downarrow$
Token
Count
$\downarrow$
Genfic EPI No 0.980 0.980 0.970 0.990 0.010 0.007 -
Embeddings No 0.955 0.955 0.95 0.960 0.040 0.049 -
Combined No 0.990 0.990 0.990 0.990 0.010 0.052 -
EPI Yes 0.980 0.980 0.960 1.000 0.000 11.041 18052.9
Embeddings Yes 0.945 0.944 0.930 0.959 0.040 12.274 18402.9
Combine Yes 0.985 0.985 0.980 0.990 0.010 10.232 18404.1
DPI Misinformation EPI No 0.610 0.625 0.650 0.602 0.430 0.008 -
Embeddings No 0.895 0.893 0.880 0.907 0.090 0.076 -
Combined No 0.680 0.698 0.740 0.661 0.380 0.086 -
EPI Yes 0.715 0.637 0.500 0.877 0.070 10.173 48256.4
Embeddings Yes 0.845 0.817 0.690 1.000 0.000 9.058 43967.9
Combined Yes 0.735 0.679 0.560 0.862 0.09 13.705 53576.6
DPI Exhaustion EPI No 0.785 0.823 1.000 0.699 0.430 0.021 -
Embeddings No 0.955 0.957 1.000 0.916 0.090 0.288 -
Combined No 0.810 0.840 1.000 0.725 0.380 0.296 -
EPI Yes 0.930 0.932 0.960 0.906 0.100 14.264 57154.2
Embeddings Yes 0.975 0.975 0.980 0.970 0.030 10.985 41082.2
Combined Yes 0.945 0.945 0.940 0.940 0.050 13.705 53576.6
DPI Backdoor EPI No 0.570 0.570 0.570 0.570 0.430 0.009 -
Embeddings No 0.890 0.888 0.870 0.906 0.090 0.081 -
Combined No 0.690 0.710 0.790 0.667 0.380 0.088 -
EPI Yes 0.740 0.680 0.550 0.887 0.070 10.927 42188.8
Embeddings Yes 0.905 0.898 0.840 0.966 0.030 10.219 41285.8
Combined Yes 0.815 0.793 0.710 0.899 0.080 12.546 50288.8
Table 16. Anomaly explanation human evaluation scores.
Type DPI IPI MP Bias Hallucination
Pattern
Detection (0–4)
2.7 2.9 2.8 2.9 2.7
Evidence
Quality (0–3)
2.2 2.3 2.3 2.5 2.3
Completeness
(0–3)
2.0 2.1 2.1 2.2 2.3
Total Score
(0–10)
6.9 7.3 7.3 7.6 7.3
C.2.13. Training Resource consumption
Table 17. Average training time (in hours) for different anomaly detection approaches using GPU and CPU